# Morpho-Photometric Classification of KiDS DR5 Sources Based on Neural Networks: A Comprehensive Star-Quasar-Galaxy Catalog

### v1

@@ Abstract
We present a novel multimodal neural network for classifying astronomical sources in multiband ground-based observations, from optical to near infrared, to separate sources in stars, galaxies and quasars. Our approach combines a convolutional neural network branch for learning morphological features from $r$-band images with an artificial neural network branch for extracting spectral energy distribution (SED) information. Specifically, we have used 9-band optical ($ugri$) and NIR ($ZYHJK_{s}$) data from the Kilo-Degree Survey (KiDS) Data Release 5. The two branches of the network are concatenated and feed into fully-connected layers for final classification. We train the network on a spectroscopically confirmed sample from the Sloan Digital Sky Survey cross-matched with KiDS. The trained model achieves 98.76% overall accuracy on an independent testing dataset, with F1 scores exceeding 95% for each class. Raising the output probability threshold, we obtain higher purity at the cost of a lower completeness. We have also validated the network using external catalogs cross-matched with KiDS, correctly classifying 99.74% of a pure star sample selected from Gaia parallaxes and proper motions, and 99.74% of an external galaxy sample from the Galaxy and Mass Assembly survey, adjusted for low-redshift contamination. We apply the trained network to 27,334,751 KiDS DR5 sources with $r\leqslant 23$ mag to generate a new classification catalog. This multimodal neural network successfully leverages both morphological and SED information to enable efficient and robust classification of stars, quasars, and galaxies in large photometric surveys.
@@ 1 Introduction
SED-based classification methods typically require a broad wavelength coverage. Depending on the classes one wants to classify these might include other bands like X-rays, infrared or radio. However, most survey projects operate within relatively narrow wavelength ranges. For instance, HSC and DES focus on optical bands, the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006) on near-infrared (NIR) bands, and the Wide-Field Infrared Survey Explorer (WISE; Wright et al., 2010) on mid-infrared bands. Classifying on the basis of photometric data points with short wavelength intervals may be affected by degeneracy (Buchs et al., 2019). One approach to address this is to combine multiple survey data (e.g., Ilbert et al., 2009; Fotopoulou et al., 2016; Khramtsov et al., 2021; Salvato et al., 2022; Yang & Shen, 2023), but this approach can only be applied to targets with overlapping regions and is limited to the depth of the shallowest survey. Traditional morphology-based classification demands high-quality data where low-surface brightness galaxies may be missed. Additionally, classifications suffer intrinsic degeneracies like compact galaxies and high-redshift QSOs confused as point sources. Combining the SED of celestial objects with their morphological features holds promise in mitigating degeneracies during the classification process and fully exploring the scientific potential of survey data (Sevilla-Noarbe et al., 2018; Khramtsov et al., 2019).
@@
Machine learning (ML) technology, with its outstanding speed and accuracy, has been widely applied in the field of astronomical big data processing (e.g., Dieleman et al., 2015; Bai et al., 2019; Logan & Fotopoulou, 2020; Szklenár et al., 2020; Chen, 2021; Zhou et al., 2021; Dubois et al., 2022; Li et al., 2022; Xie et al., 2023). Unlike traditional methods relying on physical modeling, ML can directly learn and mine the intrinsic features of celestial objects from vast observational data, sometimes even making new discoveries beyond existing theories. Currently, many ML algorithms have been successfully utilized in astronomical classification tasks. Techniques like Random Forests, extreme gradient boosting (XGBoost), support vector machines (SVM), and artificial neural networks (ANN) are commonly used for analyzing catalogs of photometric or morphological parameters, while convolutional neural networks (CNN) can directly extract information from images or spectra (Cabayol et al., 2019; Li et al., 2019; Tadaki et al., 2020; Nakazono et al., 2021; Chaini et al., 2023). Although some studies have attempted to feed both SED and morphological information into ML models (Fadely et al., 2012; Khramtsov et al., 2019), most studies still adopt parameterized morphological parameters obtained through traditional methods, which can be affected by noise.
@@
Motivated by the potential to enhance classification performance, we construct a multimodal neural network that combines the flexibility of ANN and CNN to simultaneously extract SED and image features, respectively. We have applied this network to the KiDS DR5 dataset to generate a new Star-QSO-Galaxy catalog. In Section 2, we provide a brief introduction to the data from each survey. Section 3 presents the methodology, detailing the construction and training of the multimodal neural network. The findings are discussed in Section 4, followed by the concluding remarks in Section 5.
@@ 2.1 KiDS
For each object, KiDS DR5 provides its 9-band Gaussian Aperture and PSF (GAaP; Kuijken et al., 2015) magnitudes, and then 8 adjacent band colors (e.g., $u-g$, $g-r$, etc.) can be derived. The GAaP technique is specifically designed to improve color measurement accuracy by homogenizing the PSF across different filters. This is especially important for galaxies and some nearby QSOs. To avoid any confusion introduced by the unreliable measurements, we only adopt the 65,909,027 objects which have all 9-band measurements. Figure 1 shows the distribution of their $r$-band GAaP magnitude and corresponding $u-g$ color. Considering that supervised machine learning algorithms are limited by the feature space coverage of training samples (see Figure 1), we further select the 27,334,751 objects with $r\leqslant$ 23 mag as inference data (see Section 4). In principle, the reliability of the magnitude measurement should be characterized by the error or the signal-to-noise ratio (S/N), but this can be ignored for bright sources.
@@ 2.2 SDSS
DR17 is the final annual release of SDSS-IV (Abdurro’uf et al., 2022) which has cumulated 5,801,200 spectroscopically identified objects in specObj-dr17111https://data.sdss.org/datamodel/files/SPECTRO_REDUX/specObj.html data base. We apply TOPCAT tool (Taylor, 2005) to cross-match SDSS sample with KiDS, and 177,843 objects are obtained within a matching radius of 1 arcsec. To ensure the reliability of SDSS identification, we only select the confident spectroscopic classification measurements depending on the ZWARNING flag (ZWARNING = 0; Bolton et al., 2012), which yield 167,009 objects. The final cross-matched sample includes 31,306 stars, 19,608 QSOs, and 116,095 galaxies. We then randomly split a subset of 127,009 as the training dataset to train the multimodal neural network model described in Section 3, while the remaining 40,000 sources are equally divided into validation and testing datasets. Note that the inference data are deeper than the labeled sample (see Figure 1), thus it is necessary to test the performance of our algorithm.
@@ 3 Method
We employed two standard models, ANN and CNN, both of which are categories of deep learning and significant branches of ML (Smith & Geach, 2023). These models are designed based on the concept of emulating the operations of biological neural networks, which in turn allow for the efficient processing and analysis of complex data. ANN (McCulloch et al., 1943), also referred to as Multi-Layer Perceptron (MLP), comprises multiple interconnected layers of artificial neurons that mimic the connections and synaptic processing of biological brains. Each neuron learns weighting and bias parameters to capture nonlinear relationships within inputs, thus enabling accurate prediction and classification. CNN (Lecun et al., 1998), on the other hand, is a specialized neural network architecture that is particularly effective in image and signal processing tasks. By employing components such as convolutional layers, pooling layers, and fully connected layers, CNN is capable of extracting spatial and morphological features from input data. It accomplishes this through weight sharing and local perception mechanisms. As a result, CNN exhibits superior performance in tasks like image classification and object detection, where its ability to analyze complex visual information is particularly advantageous.
Figure 2: Schematic of the multimodal neural network architecture, consisting of 18,257,155 neurons. Each blue or purple rectangle represents a layer, and each light pink rectangle represents two identical ResNet blocks. The SED branch inputs photometry from 9 KiDS bands and 8 adjacent band colors, and the image branch processes a $41\times 41$ pixels $r$-band image. The output layer provides classification probabilities for stars, quasars, or galaxies.
@@ 3.1 Neural network architecture
As illustrated in Figure 2, our multimodal neural network integrates two simultaneous branches tailored for feature extraction from SED and image data. These branches aim to capture different attributes of the data: the ANN is tasked with analyzing complex nonlinear patterns in multi-band photometric information, while the CNN focuses on extracting spatial structures and morphological features from astronomical images. The high-level features extracted from both data streams are combined to form a complementary feature vector that feeds into the final classification layers, thereby improving the overall accuracy of our model.
The ANN branch begins with an input layer, where it receives photometric information spanning nine bands along with eight colors derived from adjacent bands. Following this are four fully connected layers, each harboring 512 neurons. These dense layers are dedicated to capturing the complex non-linear relationships embedded within the SED data. The output from the final dense layer is subsequently flattened into a one-dimensional vector containing 8704 elements.
@@
The feature vectors from the ANN and CNN—8704 and 512 dimensions, respectively—are concatenated to form a comprehensive 9216-dimensional vector that incorporates both the photometric and morphological characteristics of the astronomical sources. This integrated vector is then refined through a series of dense layers with a descending number of neurons (512, 256, 128, and 64), progressively reducing the dimensionality of the feature space and concentrating on the most salient features. Culminating the network is a final dense layer with three neurons, corresponding to the categories of stars, galaxies, and QSOs. This output layer employs a softmax activation function to produce a probabilistic distribution over the three classes, facilitating the determination of the most likely category for each astronomical source.
Throughout the network, each hidden layer applies the ReLU activation function to ensure non-linearity and prevent gradient saturation, which is vital for maintaining gradient flow during the training process. We adopted the categorical cross-entropy as our loss function and utilized the Adam optimizer to refine the training of our model.
@@ 3.2 Evaluation metrics
Precision, also known as Purity in astronomy, quantifies the fraction of true positives among the predicted positives. Recall represents the proportion of all positive samples that are successfully identified by the model, equivalent to Completeness in astronomy. It is important to note that ML practitioners often enhance classification confidence by setting a threshold for output probabilities, potentially excluding data below the threshold. When calculating Recall, some studies count these below-threshold samples as FN, while others completely ignore them. In scenarios where they are included, Recall still aligns with Completeness, otherwise, Recall needs to be adjusted by multiplying the fraction of samples above the threshold ($\rm Frac_{thre}$). Therefore, in our analysis of the testing dataset, we compute $\rm Frac_{thre}$ for different thresholds. The F1-score, the harmonic average of Precision and Recall, is particularly valuable as it provides a balanced measure of classification performance, reflecting the overall efficacy of the model across each category of sources. This metric was instrumental in evaluating the ability of our multimodal neural network to classify astronomical sources. Figure 3 displays the distribution of these evaluation metrics across different output probability thresholds, with corresponding values tabulated in Table 1.
Additionally, the confusion matrix is a powerful tool for evaluating classification problems, offering a visual representation of the relationship between model predictions for each class and the actual ground-truth categories. Figure 4 illustrates the results at an output probability threshold of 0.995, where each column represents the number of predictions for the corresponding class, and each row represents the ground-truth classes. To render the classification precision of each category more discernible, we have normalized the confusion matrix and displayed it in a percentage format.
Table 1: Classification Performance at Different Thresholds.
Threshold Class Precision Recall F1-score $\rm Frac_{thre}$ Completeness
$p\geqslant$ pmax Star 98.65 98.62 98.63 100.00 98.62
QSO 96.71 94.53 95.61 100.00 94.53
Galaxy 99.09 99.44 99.26 100.00 99.44
$p\geqslant$ 0.5 Star 98.65 98.62 98.63 100.00 98.62
QSO 96.76 94.53 95.63 100.00 94.53
Galaxy 99.09 99.44 99.27 99.98 99.42
$p\geqslant$ 0.8 Star 98.91 99.16 99.03 98.48 97.65
QSO 98.19 96.30 97.24 95.64 92.10
Galaxy 99.44 99.66 99.55 99.18 98.85
$p\geqslant$ 0.9 Star 99.06 99.32 99.19 97.10 96.44
QSO 98.79 97.08 97.93 92.93 90.22
Galaxy 99.58 99.76 99.67 98.23 98.00
$p\geqslant$ 0.99 Star 99.70 99.66 99.68 82.38 82.10
QSO 99.45 98.69 99.07 84.29 83.19
Galaxy 99.82 99.94 99.88 85.74 85.69
$p\geqslant$ 0.995 Star 99.77 99.77 99.77 72.32 72.15
QSO 99.66 99.09 99.37 80.66 79.93
Galaxy 99.87 99.95 99.91 78.19 78.15
$p\geqslant$ 0.999 Star 99.79 99.79 99.79 39.23 39.14
QSO 99.73 99.66 99.69 67.34 67.11
Galaxy 99.96 99.97 99.97 52.74 52.73
Note. — Evaluation metrics including Precision, Recall, F1-score, $\rm Frac_{thre}$, and Completeness at different output probability thresholds for stars, quasars, and galaxies.
@@ 3.3 Implementation
During the training phase, the hyperparameters such as batch size and regularization factor were determined through repeated experiments. We found that setting the batch size and regularization factor to 128 and 0.001, respectively, was sufficient to achieve good performance on the testing dataset (containing 20,000 samples). To rapidly and accurately pinpoint the optimal model, we employed a tiered learning rate strategy, conducting training at learning rates of 1e-3, 1e-4, and 1e-5, across 30 epochs for each rate. Each epoch took approximately 34 seconds, with a total training time of $\sim$51 minutes. We then selected the best model based on the minimum loss and maximum accuracy on the validation dataset. This model achieved nearly identical performance on the testing dataset, and we arbitrarily chose the model with the lowest loss to make predictions on the full dataset.
Figure 5: Same as Figure 4 but for confusion matrices at different output probability thresholds, showing how the predictive performance changes with the adjustment of the threshold.
4 Results and Discussion
We assessed the overall classification capability and applicability of the proposed multimodal neural network by examining its performance on an independent internal testing dataset. Furthermore, two external datasets were utilized to validate the robustness and generalization ability of the model beyond the training sample space. Finally, we applied the trained network to the full KiDS DR5 dataset and generated a new Star-QSO-Galaxy catalog for objects brighter than 23 mag in $r$-band.

### v2

@@ Abstract
We present a novel multimodal neural network (MNN) for classifying astronomical sources in multiband ground-based observations, from optical to near infrared, to separate sources in stars, galaxies and quasars. Our approach combines a convolutional neural network branch for learning morphological features from $r$-band images with an artificial neural network branch for extracting spectral energy distribution (SED) information. Specifically, we have used 9-band optical ($ugri$) and NIR ($ZYHJK_{s}$) data from the Kilo-Degree Survey (KiDS) Data Release 5. The two branches of the network are concatenated and feed into fully-connected layers for final classification. We train the network on a spectroscopically confirmed sample from the Sloan Digital Sky Survey cross-matched with KiDS. The trained model achieves 98.76% overall accuracy on an independent testing dataset, with F1 scores exceeding 95% for each class. Raising the output probability threshold, we obtain higher purity at the cost of a lower completeness. We have also validated the network using external catalogs cross-matched with KiDS, correctly classifying 99.74% of a pure star sample selected from Gaia parallaxes and proper motions, and 99.74% of an external galaxy sample from the Galaxy and Mass Assembly survey, adjusted for low-redshift contamination. We apply the trained network to 27,335,836 KiDS DR5 sources with $r\leqslant 23$ mag to generate a new classification catalog. This MNN successfully leverages both morphological and SED information to enable efficient and robust classification of stars, quasars, and galaxies in large photometric surveys.
@@ 1 Introduction
SED-based classification methods typically require a broad wavelength coverage. Depending on the classes one wants to classify these might include other bands like X-rays, infrared or radio. However, most survey projects operate within relatively narrow wavelength ranges. For instance, HSC and DES focus on optical bands, the Two Micron All Sky Survey (2MASS; Skrutskie et al., 2006) on near-infrared (NIR) bands, and the Wide-Field Infrared Survey Explorer (WISE; Wright et al., 2010) on mid-infrared bands. Classifying on the basis of photometric data points with short wavelength intervals may be affected by degeneracy (Buchs et al., 2019). One approach to address this is to combine multiple survey data (e.g., Ilbert et al., 2009; Fotopoulou et al., 2016; Khramtsov et al., 2021; Salvato et al., 2022; Yang & Shen, 2023), but this approach can only be applied to targets with overlapping regions and is limited to the depth of the shallowest survey. Traditional morphology-based classification demands high-quality data where low-surface brightness galaxies may be missed. Additionally, classifications suffer intrinsic degeneracies like compact galaxies and high-redshift QSOs confused as point sources. Combining the SED of celestial objects with their morphological features holds promise in mitigating degeneracies during the classification process and fully exploring the scientific potential of survey data (Sevilla-Noarbe et al., 2018; Khramtsov et al., 2019; Nakoneczny et al., 2019, 2021).
@@
Machine learning (ML) technology, with its outstanding speed and accuracy, has been widely applied in the field of astronomical big data processing (e.g., Dieleman et al., 2015; Bai et al., 2019; Logan & Fotopoulou, 2020; Szklenár et al., 2020; Chen, 2021; Zhou et al., 2021; Dubois et al., 2022; Li et al., 2022; Xie et al., 2023). Unlike traditional methods relying on physical modeling, ML can directly learn and mine the intrinsic features of celestial objects from vast observational data, sometimes even making new discoveries beyond existing theories. Currently, many ML algorithms have been successfully utilized in astronomical classification tasks. Techniques like Random Forests, extreme gradient boosting (XGBoost), support vector machines, and artificial neural networks (ANN) are commonly used for analyzing catalogs of photometric or morphological parameters, while convolutional neural networks (CNN) can directly extract information from images or spectra (Cabayol et al., 2019; Li et al., 2019; Tadaki et al., 2020; Nakazono et al., 2021; Chaini et al., 2023).
Although some studies have attempted to feed both SED and morphological information into ML models (Fadely et al., 2012; Khramtsov et al., 2019; Nakoneczny et al., 2021), most studies still rely on parameterized morphological parameters obtained through traditional methods, which can be affected by random noise. Meanwhile, these methods often face the following limitations: (1) they are generally more suited for small-scale datasets, with performance plateauing as dataset size increases; (2) they struggle to model complex patterns that require deep and hierarchical representations; and (3) they require manual feature engineering, which is labor-intensive and prone to biases.
@@
Motivated by the potential to enhance classification performance, we construct an multimodal neural network (MNN) that combines the flexibility of ANN and CNN to simultaneously extract SED and image features, respectively. We have applied this network to the KiDS DR5 dataset to generate a new Star-QSO-Galaxy catalog. In Section 2, we provide a brief introduction to the data from each survey. Section 3 presents the methodology, detailing the construction and training of the MNN. The results are presented in Section 4, followed by a discussion in Section 5 and concluding remarks in Section 6.
@@ 2.1 KiDS
For each object, KiDS DR5 provides its 9-band Gaussian Aperture and PSF (GAaP; Kuijken et al., 2015) magnitudes, and then 8 adjacent band colors (e.g., $u-g$, $g-r$, etc.) can be derived. The GAaP technique is specifically designed to improve color measurement accuracy by homogenizing the PSF across different filters. This is especially important for galaxies and some nearby QSOs. To avoid any confusion introduced by the unreliable measurements, we only adopt the 65,909,027 objects which have all 9-band measurements. Figure 1 shows the distribution of their $r$-band GAaP magnitude and corresponding $u-g$ color. Considering that supervised machine learning algorithms are limited by the feature space coverage of training samples (see Figure 1), we further select the 27,335,836 objects with $r\leqslant$ 23 mag as inference data (see Section 4). In principle, the reliability of the magnitude measurement should be characterized by the error or the signal-to-noise ratio (S/N), but this can be ignored for bright sources.
@@ 2.2 SDSS
DR17 is the final annual release of SDSS-IV (Abdurro’uf et al., 2022) which has cumulated 5,801,200 spectroscopically identified objects in specObj-dr17111https://data.sdss.org/datamodel/files/SPECTRO_REDUX/specObj.html data base. We apply TOPCAT tool (Taylor, 2005) to cross-match SDSS sample with KiDS, and 177,843 objects are obtained within a matching radius of 1 arcsec. To ensure the reliability of SDSS identification, we only select the confident spectroscopic classification measurements depending on the ZWARNING flag (ZWARNING = 0; Bolton et al., 2012), which yield 167,009 objects. The final cross-matched sample includes 31,306 stars, 19,608 QSOs, and 116,095 galaxies. We then randomly split a subset of 127,009 (76%) as the training dataset to train the MNN model described in Section 3, while the remaining 40,000 (24%) sources are equally divided into validation and testing datasets. Note that the inference data are deeper than the labeled sample (see Figure 1), thus it is necessary to test the performance of our algorithm.
@@ 3 Method
We employed two standard models, ANN and CNN, which are significant branches of ML (Smith & Geach, 2023). These models are designed based on the concept of emulating the operations of biological neural networks, which in turn allow for the efficient processing and analysis of complex data. ANN (McCulloch et al., 1943), also referred to as Multi-Layer Perceptron (MLP), comprises multiple interconnected layers of artificial neurons that mimic the connections and synaptic processing of biological brains. When an ANN comprises multiple hidden layers, it belongs to the category of deep neural networks (DNNs). Each neuron learns weighting and bias parameters to capture nonlinear relationships within inputs, thus enabling accurate prediction and classification. CNN (Lecun et al., 1998), on the other hand, is a specialized neural network architecture that is inherently a part of the DNN family. It is particularly effective in image and signal processing tasks. By employing components such as convolutional layers, pooling layers, and fully connected layers, CNN is capable of extracting spatial and morphological features from input data. It accomplishes this through weight sharing and local perception mechanisms. As a result, CNN exhibits superior performance in tasks like image classification and object detection, where its ability to analyze complex visual information is particularly advantageous.
Figure 2: Schematic of the MNN architecture, consisting of 18,257,155 neurons. Each blue or purple rectangle represents a layer, and each light pink rectangle represents two identical ResNet blocks. The SED branch inputs photometry from 9 KiDS bands and 8 adjacent band colors, and the image branch processes a $41\times 41$ pixels $r$-band image. The output layer provides classification probabilities for stars, quasars, or galaxies.
@@ 3.1 Neural network architecture
As illustrated in Figure 2, our MNN integrates two simultaneous branches tailored for feature extraction from SED and image data. These branches aim to capture different attributes of the data: the ANN is tasked with analyzing complex nonlinear patterns in multi-band photometric information, while the CNN focuses on extracting spatial structures and morphological features from astronomical images. The high-level features extracted from both data streams are combined to form a complementary feature vector that feeds into the final classification layers, thereby improving the overall accuracy of our model.
The ANN branch begins with an input layer, where it receives photometric information spanning nine bands along with eight colors derived from adjacent bands. Following this are four fully connected layers, consisting of 64, 128, 256, and 512 neurons, respectively. These dense layers are dedicated to capturing the complex non-linear relationships embedded within the SED data. The output from the final dense layer is subsequently flattened into a one-dimensional vector containing 8704 elements.
@@
The architecture of the MNN was developed through a systematic and iterative process. For the classification of SEDs, we first evaluated dense networks with 1 to 15 layers and found that performance stabilized beyond 4 layers. Based on this result, a 4-layer dense architecture was adopted to construct the MLP component of the network. This MLP was then concatenated with the ResNet18 backbone to enable multimodal feature fusion for classification. During this stage, we systematically explored various CNN hyperparameters, such as the number of layers, pooling strategies, and kernel sizes, and found that some adjustments contributed to performance improvements. Additionally, we examined the addition of dense layers at the output stage, which yielded further gains. The final architecture was refined by fine-tuning parameters across the entire multimodal network to achieve optimal results. Throughout the design process, we prioritized computational efficiency by minimizing the number of parameters when performance differences were negligible. The performance differences across most network configurations were typically less than 1%.
The feature vectors from the ANN and CNN—8704 and 512 dimensions, respectively—are concatenated to form a comprehensive 9216-dimensional vector that incorporates both the photometric and morphological characteristics of the astronomical sources. This integrated vector is then refined through a series of dense layers with a descending number of neurons (512, 256, 128, and 64), effectively reducing the dimensionality of the feature space. The progressive dimensionality reduction may serve multiple purposes within the network. For example, it helps mitigate the risk of overfitting by simplifying the feature space, facilitates the integration and transformation of features from different modalities, and improves computational efficiency by reducing the overall complexity of the network (Bengio et al., 2012). Culminating the network is a final dense layer with three neurons, corresponding to the categories of stars, galaxies, and QSOs. This output layer employs a softmax activation function to produce a probabilistic distribution over the three classes, facilitating the determination of the most likely category for each astronomical source.
Throughout the network, each hidden layer applies the ReLU activation function to introduce non-linearity and facilitate effective gradient propagation, which is vital for maintaining efficient training of deep neural networks (Nair & Hinton, 2010; Glorot et al., 2011). We adopted the categorical cross-entropy as our loss function and utilized the Adam optimizer to refine the training of our model.
@@
Figure 5: Training and validation loss (top) and accuracy (bottom) as a function of epochs.
@@ 3.2 Evaluation metrics
Precision, also known as Purity in astronomy, quantifies the fraction of true positives among the predicted positives. Recall represents the proportion of all positive samples that are successfully identified by the model, equivalent to Completeness in astronomy. It is important to note that ML practitioners often enhance classification confidence by setting a threshold for output probabilities, potentially excluding data below the threshold. When calculating Recall, some studies count these below-threshold samples as FN, while others completely ignore them (Nakoneczny et al., 2021; Yang & Shen, 2023). In scenarios where they are included, Recall still aligns with Completeness, otherwise, Recall needs to be adjusted by multiplying the fraction of samples above the threshold ($\rm Frac_{thre}$). Therefore, in our analysis of the testing dataset, we compute $\rm Frac_{thre}$ for different thresholds. The F1-score, the harmonic average of Precision and Recall, is particularly valuable as it provides a balanced measure of classification performance, reflecting the overall efficacy of the model across each category of sources. This metric was instrumental in evaluating the ability of our MNN to classify astronomical sources. Figure 3 displays the distribution of these evaluation metrics across different output probability thresholds.
Additionally, the confusion matrix is a powerful tool for evaluating classification problems, offering a visual representation of the relationship between model predictions for each class and the actual ground-truth categories. Figure 4 illustrates the results at the maximum output probability ($p_{\rm max}$), where each column represents the number of predictions for the corresponding class, and each row represents the ground-truth classes. To render the classification precision of each category more discernible, we have normalized the confusion matrix and displayed it in a percentage format.
@@ 3.3 Implementation
During the training phase, the hyperparameters such as batch size and regularization factor were determined through repeated experiments. We found that setting the batch size and L2 regularization factor to 128 and 0.001, respectively, was sufficient to achieve good performance on the testing dataset (containing 20,000 samples). To efficiently pinpoint the optimal model, we tested multiple learning rates ranging from 1e-3 to 1e-5. The results revealed minimal performance differences among these rates, and we ultimately selected 1e-4 as the learning rate due to its stability and computational efficiency. Training was conducted for a total of 50 epochs, as the validation loss and accuracy were observed to plateau after approximately 20 epochs. The training and validation loss, along with the accuracy for each epoch, are presented in Figure 5. Each epoch took approximately 32 seconds, with a total training time of $\sim$27 minutes. We then selected the best model based on the minimum loss and maximum accuracy on the validation dataset. This model achieved nearly identical performance on the testing dataset, and we arbitrarily chose the model with the lowest loss to make predictions on the full dataset.
Table 1: Classification Performance for Different Methods.
Method Star QSO Galaxy
Precision Recall F1-score Precision Recall F1-score Precision Recall F1-score
MNN (Img1+Pho9) 0.9826 0.9908 0.9867 0.9630 0.9325 0.9475 0.9896 0.9926 0.9911
MNN (Img1+Pho4) 0.9558 0.9309 0.9432 0.9003 0.8838 0.8920 0.9851 0.9944 0.9897
MNN (Img4+Pho4) 0.9543 0.9337 0.9439 0.9086 0.8810 0.8946 0.9854 0.9954 0.9903
CNN (Img4) 0.8982 0.8876 0.8929 0.8163 0.7859 0.8008 0.9816 0.9901 0.9858
MPL (Pho9) 0.9709 0.9730 0.9720 0.9344 0.8757 0.9041 0.9791 0.9889 0.9840
XGBoost (Pho9) 0.9854 0.9854 0.9854 0.9534 0.9152 0.9339 0.9865 0.9916 0.9891
Note. — Evaluation metrics (Precision, Recall, and F1-score) for stars, quasars, and galaxies using different classification methods. ”Img” and ”Pho” indicate the number of imaging bands and photometric bands, respectively, used during the training process.
4 Results
We assessed the overall classification capability and applicability of the proposed MNN by examining its performance on an independent internal testing dataset. Furthermore, two external datasets were utilized to validate the robustness and generalization ability of the model beyond the training sample space. Finally, we applied the trained network to the full KiDS DR5 dataset and generated a new Star-QSO-Galaxy catalog for objects brighter than 23 mag in $r$-band.